# Simplifying FFT-based methods for mechanics with automatic differentiation

### v1

@@ 2 Differentiable FFT-framework for solid mechanics
Lastly, we exploit the automatic differentiation technique to compute the homogenized stiffness tensor, which is the most significant aspect of computational homogenization [27, 28, 29] or the essential requirement for performing multiscale simulations [12, 30]. The effective stiffness tensor is the partial derivative of macroscale stresses with respect to macroscale strains, which reads
@@
where $\overline{\bm{\sigma}}$ is the volume average of local stresses within the RVE. In the FFT-based framework, two methods are used to determine the effective properties from the local stress and strain distribution: (a) the perturbation approach [12, 31] and (b) the algorithmic consistent approach [32]. In the perturbation approach, each component of the effective tensor is computed by numerical differentiation (finite difference) between the perturbed state and the reference state. In the algorithmic consistent approach, the consistent tangent operator is computed by solving the Lippmann-Schwinger equation [32]. Both approaches necessitate an auxiliary solution method alongside the primary one to determine effective properties. For instance, the perturbation method requires computing finite differences for every tensor component, leading to three computations in a 2D case and six in a 3D case. Similarly, resolving the Lippmann-Schwinger equation requires an iterative solver. Consequently, both approaches can be computationally intensive, especially for large computational domains or for nonlinear constitutive laws. In contrast to these approaches, AD significantly simplifies the procedure by employing a single operation for the computation of the homogenized tangent stiffness, as shown below: {minted}[frame=lines, framesep=1mm, baselinestretch=0.7, fontsize=, ]python # function to solve RVE using FFT approach for a given macroscale strain def solve_microscale(macro_strain): # solve RVE (Eq 1) using Newton-Krylov solver as given in Appendix A # and get the microscale strain local_strain = netwon_krylov_solver(macro_strain)
@@
# calling tangent function to get macroscale stresses and tangents macro_tangent, macro_stress = compute_macro_tangent_stiffness(macro_strain) Here, we directly differentiate a function that takes macro-strain as input and outputs macro-stress (after solving Equation 2) to compute the homogenized stiffness tensor. This function encompasses numerous complex operations or function calls, such as forward-FFT transform, inverse FFT transform, and a Newton-Krylov solver, to accurately determine the local strain and stress distribution in an RVE (for details, refer [6, 24]). The ability of AD to decompose complicated arithmetic operations into primitive ones allows for exact differentiation. Moreover, it facilitates the simultaneous computation of stresses and tangent stiffness, thus eliminating the need for any secondary solution method.
@@ 3 Applications
Incorporating automatic differentiation in an FFT-based framework can, as described in Section 2, greatly simplify the method, which is essential for many mechanical problems. Here, we employ the AD-enhanced Fourier-Galerkin method on problems of varying complexity to demonstrate its capabilities. Specifically, we examine mechanical problems with intricate material models, bodies undergoing large-strain deformation, and multiscale simulations. We evaluate the computational performance and accuracy of the AD-enhanced FFT-based method and show how automatic differentiation simplifies the approach. We provide the Python code for all the examples at [33]. All simulations were performed as a serial job on a Linux system using a 4-core AMD EPYC™ 9654 CPU with 16 GB RAM and a clock speed of 2.4Hz.
@@ 3.1 Automatic derivation of stress and tangent operator
We present two examples to demonstrate the applicability of AD for the computation of stress and tangent operators for material laws with fundamentally different characteristics, i.e. hyperelastic and elastoplastic materials.
@@ 3.1.1 Example on hyperelastic material
With this example, we aim to show the reliance of FFT-based frameworks on explicit formulations for stress and tangent operators can be minimized [4]. We consider hyperelastic materials, where the stress-strain relationship is directly derived from the strain energy density functional. We consider the St. Venant-Kirchhoff material for which the energy density functional is given as
@@
where $\bm{E}=\bm{F}\cdot{}\bm{F}^{\mathrm{T}}-\bm{I}$ is the Green-Lagrange strain, $\bm{F}$ is the deformation gradient, $\lambda$ and $\mu$ are Lame’s parameters. As described in Section 2, we apply AD to derive the stress function (the Jacobian of the energy density) and the tangent function (the Hessian of the energy density).
Foremost, we compare the AD implementation for St. Venant-Kirchhoff material with the implementation where the stress and tangent functions are computed from the analytical expression. Similar to [6], we consider a square RVE of dimension $\ell$ (see Figure 1a) with a square stiff inclusion (of dimension $\ell/3$) at one of its corners. The RVE is subjected to a macroscopic shear loading in terms of $\overline{\bm{F}}$ which in Voigt notation is given as $[1,1,1+0.1\ell]$. We descritize the RVE into 2992 grid points. An odd number of grid points is chosen to maintain the compatibility of the deformation gradient [3]. We solve the problem in one increment using a Newton-Krylov solver with a conjugate gradient (CG) solver for solving the linearized form of Equation 1. The tolerance for the outer Newton-Raphson (NR) loop is $10^{-5}$, and the tolerance for the inner loop of the CG solver is $10^{-8}$. We compute the local distribution of the first-Piola Kirchhoff stresses within the RVE (see Figure 1b) and the residual norm ($\delta\bm{F}/\|\overline{\bm{F}}\|$) for every NR iteration (see Figure 1c). The AD implementation gives identical results to the analytical expression, as the residual norm and the iteration counts (cumulative CG iterations for every NR iteration) to reach convergence are the same. The computational times for the simulations with AD implementation and with analytical implementation are approximately the same, around 19.5 seconds. Since we express the tangent operator as a function rather than storing it as a matrix, the memory footprint of the AD-enhanced version of our implementation is roughly 1/2 of that of the analytical expression.
@@
Figure 2: Automatic computation of tangent operator for J2 plasticity. (a) The microstructure of the considered RVE. The stiff circular inclusions have a higher yielding strength than the matrix.ribution of cumulative plastic strain $\varepsilon_{p}$ and (c) local distribution of strain within the RVE. The color map is the result of AD-enhanced FFT simulations, and the dashed contour lines represent the results of the analytical implementation of the elastoplastic material.
@@ 3.1.2 Example on elastoplastic material
Automatic differentiation is applicable to material laws in which the stresses and the tangent stiffness cannot be explicitly derived from the energy density. For instance, elastoplastic materials involve stress and tangent computations that depend on yielding criteria, plastic flow rule, and a return-mapping algorithm [34]. For demonstration purposes, we consider a heterogeneous RVE with stiff circular inclusions embedded in a soft matrix (see Figure 2a). We adopt an isotropic linear hardening model in a small-strain setting with each phase exhibiting distinct hardening exponents and yield stresses. The RVE is subjected to a macroscopic biaxial loading, in terms of $\overline{\bm{\varepsilon}}$ which in Voigt notation is given as $[10^{-2}\ell,10^{-2}\ell,0]$. The RVE is discretized into $199^{2}$ grid points. The tangent stiffness operator in elastoplastic materials depends on the elastic strain, the plastic strain, the type of yield surface, and the normals to this yield surface [35, 34]. We chose J2 plasticity for which the consistent constitutive tangent is equal to the elastic tangent when the trial state is elastic and otherwise given as,
@@
$\mathbb{K}$ $=\frac{\partial\bm{\sigma}^{t}}{\partial\bm{e}^{t}}=\mathbb{C}_{e}-\frac{6G^{2}\Delta\gamma}{\text{tr}\sigma_{\text{eq}}}I_{d}+4G^{2}\left(\frac{\Delta\gamma}{\text{tr}\sigma_{\text{eq}}}-\frac{1}{3G+nH(\epsilon_{p}^{(t)}+\Delta\gamma)}\right)^{n-1}\text{tr}N\otimes\text{tr}N\leavevmode\nobreak\ .$
@@
Automatic differentiation calculates the tangent stiffness on the fly by differentiating the stresses with respect to the strains. Unlike the previous case of hyper-elastic material, where the local stiffness tangent operator is computed as a double derivative of an energy density function, here we first implement a Python function to compute the correct stress state. This is done using a return mapping algorithm [35, 34]. Then, this function is automatically differentiated with respect to the input. For details on the numerical implementation of an elastoplastic material for an AD-enhanced FFT-based method, see the pseudo-code in B.
We observe that the local distribution of the plastic strain and the elastic strain (see Figure 2b and Figure 2c) for the AD-computed local tangent stiffness operator $\mathbb{K}$ matches the strains computed with the analytical expression of $\mathbb{K}$ (shown with dashed contour lines in Figure 2). We note that the computation time for the AD implementation and the analytical implementation was approximately the same, around 0.25 seconds. We believe that as the analytical expression for calculating the tangent stiffness grows more intricate, the analytical implementation begins to slow down due to the surge in arithmetic operations. In such scenarios, AD, with its straightforward process for implementing tangent stiffness and comparable computational performance, becomes more advantageous.
@@
Through this example, we show that a local stiffness operator can also be achieved for the cases where the stresses are not a direct derivative of the energy functional. This increases the applicability of AD to various other material models where the calculation of stresses involves complex computational functions or operations.
@@
Table 1: Components of the effective elastic tensor $\mathbb{C}$ as computed numerically by automatic differentiation using FFT solvers compared against Eshelby’s solution for a stiff inclusion embedded in a soft infinite matrix.
3.2 Computational homogenization
We apply the AD-enhanced FFT-based framework to perform computational homogenization of an RVE. As noted earlier, current approaches, such as numerical differentiation or the algorithmic consistent approach, require an auxiliary computationally intensive method to compute the homogenized stiffness tensor of an RVE. Here, we present two examples to illustrate the efficacy and efficiency of AD in computing effective properties.
First, we consider the Eshelby problem as a basic example to compare the numerical result against the analytical solution. To this end, we model a periodic micro-structure with a stiff circular inclusion embedded in a soft matrix. We compare the numerically computed effective elastic tensor $\overline{\mathbb{C}}$ with Eshelby’s solution. We consider a very low volume fraction of the inclusion to match the infinite medium assumption of Eshelby, as commonly done [32]. The material parameters chosen for the matrix are $\lambda_{m}=2$ GPa and $\mu=1$ GPa, and for the stiff inclusion are $\lambda_{i}=10$ GPa and $\mu_{i}=5$ GPa. We observe that the results from the AD implementation are close to the analytical results (see Table 1). While an in-depth comparison of computational cost with the above-mentioned approaches is beyond the scope of this paper, we note that for a computational grid of 1992, the total computation time was around 4 sec.
@@ 3.2 Computational homogenization
Next, we apply the AD-enhanced computational homogenization method to an RVE (of dimension $\ell$) with intricate microstructure. We chose architected materials as our example. Architected materials are composed of two or more distinct phase materials arranged in a specific pattern that results in exceptional mechanical properties [36, 37, 38]. We examine a two-phase linear elastic architected material as illustrated in Figure 3a, comprising a soft phase and a hard phase, where the hard phase is arranged in a hexagonal pattern. We vary the phase contrast $\mathcal{X}=\mathrm{soft/hard}$ from 1 to $10^{-4}$ where a value of 1 represents a homogeneous material and value $\ll 1$ approximates a lattice metamaterial [39]. We apply a macroscopic compressive strain $\overline{\bm{\varepsilon}}$ in x and y-directions under a small-strain setting. The macroscopic strain tensor in Voigt notation is given as $[-10^{-3}\ell,-10^{-3}\ell,0]$. For the case where $\mathcal{X}=1$, i.e. a homogeneous material, the computed macro stiffness values match the analytical solution (see Figure 3c). With increasing phase contrast, however, the macroscopic stiffness reduces, as expected, and converges to values that are indicative of lattice metamaterials. This demonstrates that the AD-enhanced approach is able to handle substantial contrast in phase properties.
@@
Figure 3: Computational homogenization of an architected lattice material. (a) The architected RVE with a hard phase arranged in a hexagonal pattern within a soft phase. (b) Local distribution of shear stresses within the RVE for a phase contrast of $\mathcal{X}=0.5$ and a macroscopic strain $\overline{\bm{e}}=[-10^{-3}\ell,-10^{-3}\ell,0]$. (c) Components of the macroscale constitutive tangent stiffness as a function of phase contrast $\mathcal{X}$. The dark grey horizontal lines show the stiffness component for the homogeneous case i.e. $\mathcal{X}=1$.

### v2

@@ 3.3 Automatic computation of homogenized stiffness tensor
We exploit the automatic differentiation technique to compute the homogenized stiffness tensor, which is the most significant aspect of computational homogenization [30, 31] or the essential requirement for performing multiscale simulations [14, 32]. The effective stiffness tensor is the partial derivative of macroscale stresses with respect to macroscale strains, which reads
@@
where $\overline{\bm{\sigma}}$ is the volume average of local stresses within the RVE. In the FFT-based framework, two methods are used to determine the effective properties from the local stress and strain distribution: (a) the perturbation approach [14, 33] and (b) the algorithmic consistent approach [34]. In the perturbation approach, each component of the effective tensor is computed by numerical differentiation (finite difference) between the perturbed state and the reference state. In the algorithmic consistent approach, the consistent tangent operator is computed by solving the Lippmann-Schwinger equation [34]. Both approaches necessitate an auxiliary solution method alongside the primary one to determine effective properties. For instance, the perturbation method requires computing finite differences for every tensor component, leading to three computations in a 2D case and six in a 3D case. Similarly, resolving the Lippmann-Schwinger equation requires an iterative solver. Consequently, both approaches can be computationally intensive, especially for large computational domains or for nonlinear constitutive laws. In contrast to these approaches, automatic differentiation significantly simplifies the procedure by employing a single operation for the computation of the homogenized tangent stiffness, as shown in Section 3.3. {listing} {minted}[frame=lines, framesep=1mm, baselinestretch=0.7, fontsize=, ]python # function to solve RVE using FFT approach for a given macroscale strain def solve_microscale(macro_strain): # solve RVE (Eq 1) using Newton-Krylov solver as given in Appendix A # and get the microscale strain local_strain = netwon_krylov_solver(macro_strain)
@@
# calling tangent function to get macroscale stresses and tangents macro_tangent, macro_stress = compute_macro_tangent_stiffness(macro_strain) Automatic derivation of homogenized stiffness tensor
Here, we directly differentiate a function that takes macro-strain as input and outputs macro-stress (after solving Equation 3) to compute the homogenized stiffness tensor. This function encompasses numerous complex operations or function calls, such as forward-FFT transform, inverse FFT transform, and a Newton-Krylov solver, to accurately determine the local strain and stress distribution in an RVE (for details, refer [7, 27]). The ability of automatic differentiation to decompose complicated arithmetic operations into primitive ones allows for exact differentiation. Moreover, it facilitates the simultaneous computation of stresses and tangent stiffness, thus eliminating the need for any secondary solution method.
@@ 4 Applications
Incorporating automatic differentiation in an FFT-based framework can, as described in Section 3, greatly simplify the computation of complex derivatives, which is essential for many mechanical problems. Here, we employ the AD-enhanced Fourier-Galerkin method on problems of varying complexity to demonstrate its capabilities. Specifically, we examine mechanical problems with intricate material models, bodies undergoing large-strain deformation, and multiscale simulations. We show how automatic differentiation allows easy differentiation of functions with complex mathematical operations such as FFT transform or with logical operations such as conditional/unconditional-loops, which simplifies the entire numerical framework. We evaluate the computational performance and accuracy of the AD-enhanced FFT-based method. We provide the Python code for all the examples at [35]. All simulations were performed as a serial job on a Linux system using a 4-core AMD EPYC™ 9654 CPU with 16 GB RAM and a clock speed of 2.4Hz.
@@ 4.1 Automatic derivation of stress and tangent operator
We present two examples to demonstrate the applicability of automatic differentiation for the computation of stress and tangent operators for material laws with fundamentally different characteristics, i.e. hyperelastic and elastoplastic materials.
@@ 4.1.1 Example on hyperelastic material
With this example, we aim to show that the reliance of FFT-based frameworks on explicit formulations for stress and tangent operators can be minimized [5]. We consider hyperelastic materials, where the stresses and the tangent moduli are first- and second-order derivatives of the strain energy density functional, respectively. We consider the St. Venant-Kirchhoff material for which the energy density functional is given as
@@
where $\bm{E}=\bm{F}\cdot{}\bm{F}^{\mathrm{T}}-\bm{I}$ is the Green-Lagrange strain, $\bm{F}$ is the deformation gradient, $\lambda$ and $\mu$ are Lame’s parameters. As described in Section 3, we apply automatic differentiation to derive the stress function (the Jacobian of the energy density) and the tangent function (the Hessian of the energy density).
Foremost, we compare the AD implementation for St. Venant-Kirchhoff material with the implementation where the stress and tangent functions are computed from the analytical expression. Similar to [7], we consider a square RVE of dimension $\ell$ (see Figure 3a) with a square stiff inclusion (of dimension $\ell/3$) at one of its corners. The RVE is subjected to a macroscopic shear loading in terms of $\overline{\bm{F}}$ which in Voigt notation is given as $[1,1,1+0.1\ell]$. We discretize the RVE into 2992 grid points. An odd number of grid points is chosen to maintain the compatibility of the deformation gradient [4]. We solve the problem in one increment using a Newton-Krylov solver with a conjugate gradient (CG) solver for solving the linearized form of Equation 2. The tolerance for the outer Newton-Raphson (NR) loop is $10^{-5}$, and the tolerance for the inner loop of the CG solver is $10^{-8}$. We compute the local distribution of the first-Piola Kirchhoff stresses within the RVE (see Figure 3b) and the residual norm ($\delta\bm{F}/\|\overline{\bm{F}}\|$) for every NR iteration (see Figure 3c). Here, we use JAX’s jvp function (see Section 3.2) to compute the incremental stresses within the CG iterations. The AD implementation gives identical results to the analytical expression, as the residual norm and the iteration counts (cumulative CG iterations for every NR iteration) to reach convergence are the same. The computational times for the simulations with AD implementation and with analytical implementation are approximately the same, around 19.5 seconds. Since we express the tangent operator as a function rather than storing it as a matrix, the memory footprint of the AD-enhanced version of our implementation is roughly 1/2 of that of the analytical expression.
@@
Figure 4: Automatic computation of tangent operator for $J_{2}$-plasticity. (a) The microstructure of the considered RVE. The hard circular inclusions have a higher yielding strength than the matrix.ribution of cumulative plastic strain $\varepsilon_{p}$ and (c) local distribution of strain within the RVE. The color map is the result of AD-enhanced FFT simulations, and the dashed contour lines represent the results of the analytical implementation of the elastoplastic material.
@@ 4.1.2 Example on elastoplastic material
Automatic differentiation is applicable to material laws in which the stresses and the tangent stiffness cannot be explicitly derived from the energy density. For instance, elastoplastic materials involve stress and tangent computations that depend on yielding criteria, plastic flow rule, and a return-mapping algorithm [36]. For demonstration purposes, we consider a heterogeneous RVE with hard circular inclusions embedded in a soft matrix (see Figure 4a). We adopt an isotropic linear hardening model in a small-strain setting with each phase exhibiting distinct hardening exponents and yield stresses. The RVE is subjected to a macroscopic biaxial loading, in terms of $\overline{\bm{\varepsilon}}$ which in Voigt notation is given as $[10^{-2}\ell,10^{-2}\ell,0]$. The RVE is discretized into $199^{2}$ grid points. The tangent stiffness operator in elastoplastic materials depends on the elastic strain, the plastic strain, the type of yield surface, and the normals to this yield surface [37, 36]. We selected $J_{2}$-plasticity for both materials. We assume a linear isotropic hardening law with the yield criterion given as $\Phi=\sqrt{3/2\bm{\sigma}_{\text{dev}}:\bm{\sigma}_{\text{dev}}}-\sigma_{y}$ where $\bm{\sigma}_{\text{dev}}$ is the deviatoric stress and $\sigma_{y}$ is the yielding stress. The phase contrast between the yield stress and hardening moduli of the two phases is 2. The analytical expression for the consistent constitutive tangent is equal to the elastic tangent when the trial state is elastic and otherwise given as
@@
$\mathbb{K}$ $=\frac{\partial\bm{\sigma}^{t+\Delta t}}{\partial\bm{e}^{t+\Delta t}}=\mathbb{C}_{\text{elas}}-\frac{6\mu^{2}\Delta\epsilon_{p}}{\text{tr}\sigma_{\text{eq}}}I_{d}+4\mu^{2}\left(\frac{\Delta\epsilon_{p}}{\text{tr}\sigma_{\text{eq}}}-\frac{1}{3\mu+H}\right)\text{tr}N\otimes\text{tr}N\leavevmode\nobreak\ .$
@@
where $\mathbb{C}_{\text{elas}}$ is the elastic stiffness matrix, $\mu$ is the shear modulus, $\epsilon_{p}$ is the cumulative plastic strain, $H$ is the hardening modulus, $\sigma_{y}$ is the yielding stress and $N$ is the direction of the deviatoric strain. Automatic differentiation calculates the tangent stiffness on the fly by differentiating the stresses with respect to the strains. Unlike the previous case of hyper-elastic material, where the local stiffness tangent operator is computed as a double derivative of an energy density function, here we first implement a Python function to compute the correct stress state. The stress state is calculated using the elastic predictor-corrector approach (for details, see B), and a return mapping algorithm [37, 36]. Then, this function is automatically differentiated with respect to the input. Here, within the NR loop, we use JAX’s linearize function (see Section 3) to create the push-forward map of tangent stiffness at a fixed strain value and then use it to calculate incremental stresses within the CG iterations. For details on the numerical implementation of an elastoplastic material for an AD-enhanced FFT-based method, see the pseudo-code in B.
@@
We observe that the local distribution of the plastic strain and the elastic strain (see Figure 4b and Figure 4c) for the AD-computed local tangent stiffness operator $\mathbb{K}$ matches the strains computed with the analytical expression of $\mathbb{K}$ (shown with dashed contour lines in Figure 4). We note that the computation time for the AD implementation and the analytical implementation was approximately the same, around 0.25 seconds. Through this example, we show that a local stiffness operator can also be achieved for the cases where the stresses are not a direct derivative of the energy functional. This increases the applicability of automatic differentiation to various other material models where the calculation of stresses involves complex computational functions or operations.
Figure 5: Computational efficiency of AD-enhanced FFT method. (a) Scaling of computational time with increase in the number of grid points for two different constitutive laws. Computations with automatic differentiation and analytical expression are shown in red and blue, respectively. For hyperelastic material, we use JAX’s jvp function to compute the incremental stresses and for elastoplastic material, we use JAX’s linearize function. (b) Memory consumption for the two material laws for the case with 5112 grid points.
4.1.3 Computational performance
We present scalability and memory usage (see Figure 5) for the two cases considered above, i.e. hyperelastic and elastoplastic. We report the total computational time required for all stages, starting with the tangent stiffness function derivation and proceeding through Newton-Krylov iterations, up to convergence. For the hyperelastic case, we applied the push-forward map, utilizing the jvp function, to calculate incremental stresses within each iteration of the Conjugate Gradient solver. As discussed in Section 3.2, constructing the push-forward map in an inner loop incurs overhead, making the AD-enhanced FFT slightly more expensive than the conventional FFT implementation with analytic expressions. Our scaling analysis shows that even with a high number of pixels, the difference in computation time remains minimal (see Figure 5a). However, memory consumption is reduced by approximately half since the tangent moduli are not stored (see Figure 5b).
In the elastoplastic case, we computed and stored the push-forward map during Newton-Raphson iterations using the linearize function. The computation time here matches the standard method using analytic expressions because automatic differentiation overhead occurs only in the Newton-Raphson steps (see Figure 5a). However, unlike before, the AD-enhanced method uses more memory compared to standard FFT, probably due to the need to store the linearized function (see Figure 5b).
Finally, we note that while we showed in Figure 2b that analytic computations cost more than the push-forward map, these analytic computations become cheaper than linearize and jvp functions for repeated executions. This occurs because the analytic expression, implemented in Numpy, is optimized for multiple uses, whereas the slight overhead of using automatic differentiation makes linearize or jvp marginally computationally more expensive.
4.2 Computational homogenization
We apply the AD-enhanced FFT-based framework to perform computational homogenization of an RVE. As noted earlier, current approaches, such as numerical differentiation or the algorithmic consistent approach, require an auxiliary computationally intensive method to compute the homogenized stiffness tensor of an RVE. Here, we present two examples to illustrate the efficacy and efficiency of automatic differentiation in computing effective properties.
4.2.1 Validation with Eshelby’s problem
[…]
@@ 4.2.1 Validation with Eshelby’s problem
Table 1: Components of the effective elastic tensor $\overline{\mathbb{C}}$ as computed numerically by automatic differentiation using FFT solvers compared against Eshelby’s solution for a stiff inclusion embedded in a compliant infinite matrix.
4.2.2 Example on architected materials
@@
We apply the AD-enhanced computational homogenization method to an RVE (of dimension $\ell$) with intricate microstructure. We chose architected materials as our example. Architected materials are composed of two or more distinct phase materials arranged in a specific pattern that results in exceptional mechanical properties [38, 39, 40]. We examine a two-phase linear elastic architected material as illustrated in Figure 6a, comprising a compliant phase and a stiff phase, where the stiff phase is arranged in a hexagonal pattern. We vary the phase contrast $\mathcal{X}=\mathrm{compliant/stiff}$ from 1 to $10^{-4}$ where a value of 1 represents a homogeneous material and value $\ll 1$ approximates a lattice metamaterial [41]. We apply a macroscopic compressive strain $\overline{\bm{\varepsilon}}$ in x and y-directions under a small-strain setting. The macroscopic strain tensor in Voigt notation is given as $[-10^{-3}\ell,-10^{-3}\ell,0]$. For the case where $\mathcal{X}=1$, i.e. a homogeneous material, the computed macro stiffness values match the analytical solution (see Figure 6c). With increasing phase contrast, however, the macroscopic stiffness reduces, as expected, and the obtained values (¯$\overline{\mathbb{C}}_{xxxx}=0.51,\overline{\mathbb{C}}_{yyyy}=0.42$ and ¯$\overline{\mathbb{C}}_{xyxy}=0.059$) saturates to the values that are indicative of a lattice metamaterials. The values obtained are similar to the values (¯$\overline{\mathbb{C}}_{xxxx}=0.57,\overline{\mathbb{C}}_{yyyy}=0.52$ and ¯$\overline{\mathbb{C}}_{xyxy}=0.15$) obtained from a full-field Finite Element homogenization under affine displacement boundary conditions (for details, please refer to C). The properties obtained under periodic boundary conditions are expected to be lower than the properties obtained under affine displacement boundary conditions [42].
While not depicted in Figure 6 we calculated the homogenized properties for a phase contrast of $10^{-10}$ to approximate zero stiffness in the compliant phase to machine precision. The obtained homogenized values remain consistent, illustrating that the AD-enhanced approach is not affected by significant contrast (approaching zero) in phase properties.
@@
Figure 6: Computational homogenization of an architected lattice material. (a) The architected RVE with a stiff phase arranged in a hexagonal pattern within a compliant phase. (b) Local distribution of shear stresses within the RVE for a phase contrast of $\mathcal{X}=0.5$ and a macroscopic strain $\overline{\bm{e}}=[-10^{-3}\ell,-10^{-3}\ell,0]$. (c) Components of the macroscale constitutive tangent stiffness as a function of phase contrast $\mathcal{X}$. The dark grey horizontal lines show the stiffness component for the homogeneous case i.e. $\mathcal{X}=1$.